OpenStudy (anonymous):

The sum of five different positive integers is 100. If the smallest of these integers is 10, what is the larges possible value of one of the other four integers?

jimthompson5910 (jim_thompson5910):

The smallest integer is 10. So the other four numbers add to 100 - 10 = 90

OpenStudy (anonymous):

Is it that simple???

jimthompson5910 (jim_thompson5910):

If x, y, z and w are the four numbers, then x+y+z+w = 90

jimthompson5910 (jim_thompson5910):

Now make x as large as possible

jimthompson5910 (jim_thompson5910):

the answer isn't 90, but it helps get you the answer

OpenStudy (anonymous):

How do you find it when you don't even know one of the letters?

jimthompson5910 (jim_thompson5910):

Let's make x as large as possible

jimthompson5910 (jim_thompson5910):

Can x = 90?

OpenStudy (anonymous):

Yes, right?

jimthompson5910 (jim_thompson5910):

but that would make y, z, and w all be 0

jimthompson5910 (jim_thompson5910):

x+y+z+w = 90 90+0+0+0 = 90

OpenStudy (anonymous):

So 84?

jimthompson5910 (jim_thompson5910):

good, x+y+z+w = 90 84 + 1 + 2 + 3 = 90

jimthompson5910 (jim_thompson5910):

anything larger for x would make y, z, or w have to become smaller (and either not be positive, or not be different)

OpenStudy (anonymous):

Thank you!! I appreciate it a lot

jimthompson5910 (jim_thompson5910):

you're welcome